# Resumming Post-Minkowskian and Post-Newtonian gravitational waveform expansions

### v2

@@ II The confluent Heun equation
a non-trivial function of $u$ to be determined. We will look for solutions in the domain $z\in[0,1]$. The general $G(z)=\sum_{\alpha}c_{\alpha}G_{\alpha}(z)$ behaves near the boundaries as
$G(z)$ $\underset{z\to 0}{\approx}$ $B_{-}(1+\ldots)+B_{+}z^{1-2\kappa_{1}}e^{x\over z}(1+\ldots)$
$G(z)$ $\underset{z\to 1}{\approx}$ $D_{-}(1+\ldots)+D_{+}(1-z)^{1-\kappa_{2}-\kappa_{3}}(1+\ldots)\,\,\,\,\,\,\,\,\,$ (10)
@@
where the dots stand for subleading terms in these limits. Specifying boundary conditions corresponds to fix two linear combinations made out of the four coefficients $B_{\pm}$, $D_{\pm}$.
We notice that the two limits (7) overlap in the region where both $z$ and $y$ are small. We refer to this region as the near zone and here we look for solutions as a double series
${\rm Near\leavevmode\nobreak\ zone}:G_{\alpha}(z)=z^{\alpha\,a{-}\kappa_{1}}\sum_{i,j=0}^{\infty}G_{\alpha ij}(a)\,z^{i}\,({\textstyle\frac{x}{z}})^{j}$ (11)
with $x\ll z\ll 1$. $a(u)$ given in (9). The quantity $a$ plays the role of a ‘Floquet exponent’, specifying the monodromy of $G_{\alpha}$ around a contour enclosing the points 0 and $x$. As we will see, the solutions of (4) and the connection formulae take a simpler form if written in terms of the variable $a$ rather than $u$. So we write
$u$ $=$ $a^{2}+\sum_{i=1}^{\infty}u_{i}(a)x^{i}$ (12)
The coefficients $G_{\alpha ij}(a)$ and $u_{i}(a)$ can be determined recursively order by order in the variable $x$ and $a(u)$ follows by inverting the series (12).
II.1 PM approximation: exterior region
@@
$c_{10}$ $=$ $u_{1}=\frac{\left[2a^{2}\left(\kappa_{1}{+}\kappa_{2}{+}\kappa_{3}{-}\frac{1}{2}\right){+}2\kappa_{1}\kappa_{2}\kappa_{3}{-}\kappa_{1}\kappa_{2}{-}\kappa_{1}\kappa_{3}{-}\kappa_{2}\kappa_{3}\right]}{4a^{2}{-}1}$
$\widehat{c}_{10}$ $=$ $-\partial_{\kappa_{1}}u_{1}=-\frac{\left(2a^{2}+2\kappa_{2}\kappa_{3}-\kappa_{2}-\kappa_{3}\right)}{4a^{2}-1}$ (16)
@@
$\gamma_{\kappa}=1+\sum_{i=1}^{\infty}\widehat{c}_{i,i-1}x^{i}={\rm exp}\left\{-\sum_{i=1}^{\infty}{\partial_{\kappa_{1}}u_{i}\over i}x^{i}\right\}\quad$ (17)
@@
The subscript ”inst” reminds the reader that this quantity is interpreted as the instanton prepotential in the gauge theory description of the Heun equation. Indeed in the limit $z\to 0$, where one of the two gauge couplings is turned off, the quiver theory reduces to a ${\cal N}=2$ supersymmetric $SU(2)$ gauge theory with three massive fundamental flavours with prepotential ${\cal F}_{\rm inst}$.
@@
$\widetilde{H}^{0}_{\alpha}(y)=y^{{\alpha+1\over 2}(2\kappa_{1}-1)}e^{y(1+\alpha)\over 2}{}_{2}F_{0}\left({\textstyle\frac{1{+}\alpha}{2}}{-}\alpha\kappa_{1}{-}a,{\textstyle\frac{1{+}\alpha}{2}}{-}\alpha\kappa_{1}{+}a;\alpha y^{-1}\right)$ (20)
@@
$B_{\alpha\alpha^{\prime}}=e^{\frac{{\rm i}\pi(1{-}\alpha^{\prime})}{2}(\kappa_{1}{-}\alpha a)}\frac{\Gamma(1{-}2\alpha a)}{\Gamma\left({\textstyle\frac{1-\alpha^{\prime}}{2}}{-}\alpha a{+}\alpha^{\prime}\kappa_{1}\right)}$ (21)
@@
II.2 PN approximation: interior region
@@
$d_{01}=\frac{2\left(a^{2}-\kappa_{1}^{2}\right)}{4a^{2}-1},\quad\widehat{d}_{11}=-\widehat{d}_{01}=\frac{\left(2\kappa_{1}-1\right)}{4a^{2}-1}$ (27)
@@
$\widetilde{H}^{1}_{\alpha}(z)=z^{a-\kappa_{1}}(1{-}z)^{{1{+}\alpha\over 2}(1{-}\kappa_{2}{-}\kappa_{3})}\,_{2}F_{1}\left({\textstyle\frac{1{+}\alpha}{2}}{-}\alpha\kappa_{2}{+}a,{\textstyle\frac{1{+}\alpha}{2}}{-}\alpha\kappa_{3}{+}a;1{+}\alpha\left(1{-}\kappa_{2}{-}\kappa_{3}\right);1{-}z\right)$ (29)
@@
$F_{\alpha\alpha^{\prime}}=\frac{\Gamma(1{+}2\alpha a)\Gamma(\alpha^{\prime}[\kappa_{2}{+}\kappa_{3}{-}1])}{\Gamma\left({\textstyle\frac{1-\alpha^{\prime}}{2}}{+}\alpha a{+}\alpha^{\prime}\kappa_{2}\right)\Gamma\left({\textstyle\frac{1-\alpha^{\prime}}{2}}{+}\alpha a{+}\alpha^{\prime}\kappa_{3}\right)}$ (30)
@@
II.3 Boundary conditions and connection formulae
@@
$G(z)$ $=$ $\sum_{\alpha}c_{\alpha}g_{\alpha}(x)G^{0}_{\alpha}({\textstyle\frac{x}{z}})+\ldots=\sum_{\alpha,\alpha^{\prime}}c_{\alpha}g_{\alpha}(x)B_{\alpha\alpha^{\prime}}\widetilde{G}^{0}_{\alpha^{\prime}}({\textstyle\frac{x}{z}})+\ldots$ (35)
@@
$G(z)$ $\underset{z\to 0}{\approx}$ $\sum_{\alpha,\alpha^{\prime}}c_{\alpha}g_{\alpha}(x)B_{\alpha\alpha^{\prime}}({\textstyle\frac{x}{z}})^{{1+\alpha^{\prime}\over 2}(2\kappa_{1}-1)}e^{x(1+\alpha^{\prime})\over 2z}e^{\frac{1+\alpha^{\prime}}{2}\partial_{\kappa_{1}}{\cal F}_{\rm inst}}$
$G(z)$ $\underset{z\to 1}{\approx}$ $\sum_{\alpha,\alpha^{\prime}}c_{\alpha}F_{\alpha\alpha^{\prime}}(1{-}z)^{{1{+}\alpha^{\prime}\over 2}(1{-}\kappa_{2}{-}\kappa_{3})}\,h_{\alpha^{\prime}}$ (36)
@@
$h_{\alpha^{\prime}}=P_{1}(1)\,{+}{1{+}\alpha^{\prime}\over 2}(1{-}\kappa_{2}{-}\kappa_{3})\widehat{P}_{1}{}^{\prime}(1)\,$ (37)
@@
$\frac{h_{+}}{h_{-}}=e^{-\partial_{\kappa_{2}+\kappa_{3}}{\cal F}_{\rm inst}}$ (38)
@@
$B_{\alpha^{\prime}}$ $=$ $\sum_{\alpha}c_{\alpha}g_{\alpha}(x)B_{\alpha\alpha^{\prime}}x^{{1+\alpha^{\prime}\over 2}(2\kappa_{1}-1)}e^{\frac{1+\alpha^{\prime}}{2}\partial_{\kappa_{1}}{\cal F}_{\rm inst}}$
@@
III Black hole perturbation
@@
III.1 Heun gravity correspondence
@@
with $\Delta=r(r-2M)$. We are interested in the $\psi_{4}$ fundamental mode [55, 8] corresponding to $s=-2$. (40) can be put into the CHE form (4) after the identifications (look at Appendix A for details)
$R(z)$ $=$ $e^{-{x\over 2z}}(1-z)^{1+\kappa_{2}+\kappa_{3}\over 2}({\textstyle\frac{x}{z}})^{{3\over 2}-\kappa_{1}}G(z)\qquad,\quad z={2M\over r}\qquad,\qquad x=4iM\omega$ (41)
$\kappa_{1}$ $=$ $\kappa_{2}^{*}={\textstyle\frac{5}{2}}+2{\rm i}M\omega\quad,\quad\kappa_{3}={\textstyle\frac{1}{2}}-2{\rm i}M\omega\quad,\quad u=(\ell+{\textstyle\frac{1}{2}})^{2}+10{\rm i}M\omega-4\omega^{2}M^{2}$
@@
$R_{\rm in}(z)$$\approx$ $\left\{\begin{array}[]{lll}B^{\rm in}_{-}e^{-{x\over 2z}}z^{\kappa_{1}-{3\over 2}}+B^{\rm in}_{+}e^{{x\over 2z}}z^{-\kappa_{1}-{1\over 2}}&\nobreak\leavevmode\nobreak\leavevmode\nobreak\leavevmode\nobreak\leavevmode\nobreak\leavevmode\nobreak\leavevmode\nobreak\leavevmode&\quad z\to 0\\ D^{\rm in}_{-}(1-z)^{1+\kappa_{2}+\kappa_{3}\over 2}&&\quad z\to 1\\ \end{array}\right.$ (44)
$R_{\rm up}(z)$$\approx$ $\left\{\begin{array}[]{lll}B^{\rm up}_{+}e^{{x\over 2z}}z^{-\kappa_{1}-{1\over 2}}&\nobreak\leavevmode\nobreak\leavevmode\nobreak\leavevmode\nobreak\leavevmode\nobreak\leavevmode\nobreak\leavevmode\nobreak\leavevmode&\quad z\to 0\\ D^{\rm up}_{-}(1-z)^{1+\kappa_{2}+\kappa_{3}\over 2}+D^{\rm up}_{+}(1-z)^{3-\kappa_{2}-\kappa_{3}\over 2}&&\quad z\to 1\\ \end{array}\right.$ (47)
@@
$G(r,r^{\prime})={1\over W}\left\{\begin{array}[]{ccc}R_{\rm in}(r^{\prime})R_{\rm up}(r)&\nobreak\leavevmode\nobreak\leavevmode\nobreak\leavevmode\nobreak\leavevmode\nobreak\leavevmode\nobreak\leavevmode\nobreak\leavevmode\hfil&r^{\prime}<r\\ R_{\rm in}(r)R_{\rm up}(r^{\prime})&\nobreak\leavevmode\nobreak\leavevmode\nobreak\leavevmode\nobreak\leavevmode\nobreak\leavevmode\nobreak\leavevmode\nobreak\leavevmode\hfil&r<r^{\prime}\\ \end{array}\right.$ (48)
@@
$R_{\alpha}(y)=e^{-{y\over 2}}(1-{\textstyle\frac{x}{y}})^{2{-}{x\over 2}}y^{{-}1{-}{x\over 2}}G^{0}_{\alpha}(y)$ (52)
@@
$B^{\rm in,up}_{\alpha^{\prime}}$ $=$ $\sum_{\alpha}c^{\rm in,up}_{\alpha}g_{\alpha}(x)B_{\alpha\alpha^{\prime}}x^{\alpha^{\prime}\kappa_{1}+1-{\alpha^{\prime}\over 2}}e^{\frac{1+\alpha^{\prime}}{2}\partial_{\kappa_{1}}{\cal F}_{\rm inst}}$
$D^{\rm in,up}_{\alpha^{\prime}}$ $=$ $\sum_{\alpha}c^{\rm in,up}_{\alpha}F_{\alpha\alpha^{\prime}}h_{\alpha^{\prime}}e^{-{x\over 2}}\,x^{{3\over 2}-\kappa_{1}}$ (54)
@@ III.2 The incoming solution
$\ {c^{\rm in}_{+}\over c^{\rm in}_{-}}=-{F_{-+}\over F_{++}}=\frac{\Gamma({-}2a)\Gamma\left(\kappa_{2}{+}a\right)\Gamma\left(\kappa_{3}{+}a\right)}{\Gamma(2a)\Gamma\left(\kappa_{2}{-}a\right)\Gamma\left(\kappa_{3}{-}a\right)}$ (59)
@@
$R_{\rm in}(y)\underset{y\to\infty}{\approx}\sum_{\alpha,\alpha^{\prime}}c^{\rm in}_{\alpha}g_{\alpha}B_{\alpha\alpha^{\prime}}e^{\alpha^{\prime}y\over 2}y^{\alpha^{\prime}\kappa_{1}+1-{\alpha^{\prime}\over 2}}e^{\frac{1+\alpha^{\prime}}{2}\partial_{\kappa_{1}}{\cal F}_{\rm inst}}$ (60)
@@
$B^{\rm in}_{\alpha^{\prime}}=\sum_{\alpha}c^{\rm in}_{\alpha}g_{\alpha}B_{\alpha\alpha^{\prime}}x^{\alpha^{\prime}\kappa_{1}+1-{\alpha^{\prime}\over 2}}e^{\frac{1+\alpha^{\prime}}{2}\partial_{\kappa_{1}}{\cal F}_{\rm inst}}$ (61)
@@
III.3 The upgoing solution
Upgoing boundary conditions correspond to setting $B^{\rm in}_{-}=0$, that according to (54) leads to

### v3

@@ 2.1 Confluent Heun-Gauge dictionary
For concreteness, we will focus on the domain $z\in[0,1]$. The general solution $G_{1}(z)=\sum_{\alpha}c_{\alpha}G^{1}_{\alpha}(z)$ behaves near the boundaries as
$G_{1}(z)$ $\underset{z\to 0}{\approx}$ $B_{-}(1+\ldots)+B_{+}z^{2m_{3}}e^{x\over z}(1+\ldots)$
$G_{1}(z)$ $\underset{z\to 1}{\approx}$ $D_{-}(1+\ldots)+D_{+}(1-z)^{m_{1}+m_{2}}(1+\ldots)\,\,\,\,\,\,\,\,\,$ (16)
@@
where the dots stand for subleading terms in these limits. Specifying boundary conditions corresponds to fix two linear combinations made out of the four coefficients $B_{\pm}$, $D_{\pm}$. We will derive connection formulae relating the coefficients $B_{\pm}$ to $D_{\pm}$ and viceversa.
2.2 PM approximation: exterior region
@@
$c_{10}$ $=$ $u_{1}=\frac{1}{2}(1-m_{1}-m_{2}-m_{3})-{2m_{1}m_{2}m_{3}\over 4a^{2}-1}$
$\widehat{c}_{10}$ $=$ $\partial_{m_{3}}u_{1}=-\frac{1}{2}-{2m_{1}m_{2}\over 4a^{2}-1}$ (20)
@@
$\gamma_{m}=1+\sum_{i=1}^{\infty}\widehat{c}_{i,i-1}x^{i}={\rm exp}\left\{\sum_{i=1}^{\infty}{\partial_{m_{3}}u_{i}\over i}x^{i}\right\}\quad$ (21)
@@
The subscript "inst" reminds the reader that this quantity is interpreted as the instanton prepotential in the gauge theory description of the Heun equation. Indeed in the limit $z\to 0$, where one of the two gauge couplings is turned off, the quiver theory reduces to a ${\cal N}=2$ supersymmetric SU(2) gauge theory with three massive fundamental flavours with prepotential ${\cal F}_{\rm inst}$.
@@
$\widetilde{H}^{0}_{\alpha}(y)=y^{-m_{3}(1+\alpha)}e^{y(1+\alpha)\over 2}{}_{2}F_{0}\left(\frac{1}{2}-a+\alpha m_{3},\frac{1}{2}+a+\alpha m_{3},\frac{\alpha}{y}\right)$ (24)
@@
$B_{\alpha\alpha^{\prime}}=e^{\frac{{\rm i}\pi}{2}\left(1-\alpha^{\prime}\right)\left({1\over 2}-\alpha a-m_{3}\right)}\frac{\Gamma(1-2\alpha a)}{\Gamma\left(\frac{1}{2}-\alpha a-\alpha^{\prime}m_{3}\right)}$ (25)
@@
2.3 PN approximation: interior region
@@
$d_{01}=\frac{1}{2}+\frac{2m_{3}\left(1-m_{3}\right)}{4a^{2}-1},\quad\widehat{d}_{01}=-\widehat{d}_{11}=\frac{2m_{3}}{4a^{2}-1}$ (31)
@@
$\widetilde{H}^{1}_{\alpha}(z)=z^{a{+}m_{3}{-}\frac{1}{2}}(1{-}z)^{\frac{1}{2}\left(m_{1}{+}m_{2}\right)\left(1{+}\alpha\right)}\,_{2}F_{1}\left(\frac{1}{2}{+}a{+}\alpha m_{1},\frac{1}{2}{+}a{+}\alpha m_{2};1{+}\alpha\left(m_{1}{+}m_{2}\right);1{-}z\right)$ (33)
@@
$F_{\alpha\alpha^{\prime}}=\frac{\Gamma(1{+}2\alpha a)\Gamma\left(-\alpha^{\prime}\left(m_{1}+m_{2}\right)\right)}{\Gamma\left({\textstyle\frac{1}{2}}+\alpha a-\alpha^{\prime}m_{1}\right)\Gamma\left({\textstyle\frac{1}{2}}+\alpha a-\alpha^{\prime}m_{2}\right)}$ (34)
@@
2.4 Boundary conditions and connection formulae
@@
$g_{\alpha}(x)={G^{1}_{\alpha}(z)\over G^{0}_{\alpha}({\textstyle\frac{x}{z}})}=x^{\alpha a{+}m_{3}{-}\frac{1}{2}}\left[1{+}x\left({\textstyle\frac{1}{4}}{-}{\textstyle\frac{m_{3}\left(m_{1}{+}m_{2}{+}2m_{3}{-}2\right){-}m_{1}m_{2}}{4a^{2}-1}}{+}{\textstyle\frac{8\,\alpha\,a\,m_{1}m_{2}m_{3}}{\left(4a^{2}{-}1\right)^{2}}}\right)+\ldots\right]$ (37)
@@
$G_{1}(z)$ $=$ $\sum_{\alpha}c_{\alpha}g_{\alpha}(x)G^{0}_{\alpha}({\textstyle\frac{x}{z}})+\ldots=\sum_{\alpha,\alpha^{\prime}}c_{\alpha}g_{\alpha}(x)B_{\alpha\alpha^{\prime}}\widetilde{G}^{0}_{\alpha^{\prime}}({\textstyle\frac{x}{z}})+\ldots$ (39)
@@
$G_{1}(z)$ $\underset{z\to 0}{\approx}$ $\sum_{\alpha,\alpha^{\prime}}c_{\alpha}g_{\alpha}(x)B_{\alpha\alpha^{\prime}}({\textstyle\frac{x}{z}})^{-m_{3}(1+\alpha^{\prime})}e^{x(1+\alpha^{\prime})\over 2z}e^{-\frac{1+\alpha^{\prime}}{2}\partial_{m_{3}}{\cal F}_{\rm inst}}$
$G_{1}(z)$ $\underset{z\to 1}{\approx}$ $\sum_{\alpha,\alpha^{\prime}}c_{\alpha}F_{\alpha\alpha^{\prime}}(1{-}z)^{{1{+}\alpha^{\prime}\over 2}(m_{1}{+}m_{2})}\,h_{\alpha^{\prime}}$ (40)
@@
$h_{\alpha^{\prime}}=P_{1}(1)\,{+}{1{+}\alpha^{\prime}\over 2}(m_{1}{+}m_{2})\widehat{P}_{1}{}^{\prime}(1)\,$ (41)
@@
$\frac{h_{+}}{h_{-}}=e^{\partial_{m_{1}+m_{2}}{\cal F}_{\rm inst}}$ (42)
@@
$B_{\alpha^{\prime}}$ $=$ $\sum_{\alpha}c_{\alpha}g_{\alpha}(x)B_{\alpha\alpha^{\prime}}x^{-m_{3}(1+\alpha^{\prime})}e^{-\frac{1+\alpha^{\prime}}{2}\partial_{m_{3}}{\cal F}_{\rm inst}}$
@@
3 Black hole perturbation
@@
3.1 Heun gravity correspondence
@@
with $\Delta=r(r-2M)$. We are interested in the $\psi_{4}$ fundamental mode Newman:1961qr ; Teukolsky:1973ha corresponding to $s=-2$. (44) can be put into the CHE form (8) after the following identifications
$R(z)$ $=$ $e^{-{x\over 2z}}(1-z)^{1-{m_{1}+m_{2}\over 2}}({\textstyle\frac{x}{z}})^{1+m_{3}}G_{1}(z)\qquad,\quad z={2M\over r}\qquad,\qquad x=4iM\omega$ (45)
$m_{1}$ $=$ $2{\rm i}M\omega\quad,\quad m_{2}=m_{3}^{*}=-2+2{\rm i}M\omega\quad,\quad u=(\ell+{\textstyle\frac{1}{2}})^{2}+10{\rm i}M\omega-4\omega^{2}M^{2}$
@@
$R_{\rm in}(z)$$\approx$ $\left\{\begin{array}[]{lll}B^{\rm in}_{-}e^{-{x\over 2z}}z^{-1-m_{3}}+B^{\rm in}_{+}e^{{x\over 2z}}z^{m_{3}-1}&{}{}{}{}{}{}{}&\quad z\to 0\\ D^{\rm in}_{-}(1-z)^{1-{m_{1}+m_{2}\over 2}}&&\quad z\to 1\\ \end{array}\right.$ (48)
$R_{\rm up}(z)$$\approx$ $\left\{\begin{array}[]{lll}B^{\rm up}_{+}e^{{x\over 2z}}z^{m_{3}-1}&{}{}{}{}{}{}{}&\quad z\to 0\\ D^{\rm up}_{-}(1-z)^{1-{m_{1}+m_{2}\over 2}}+D^{\rm up}_{+}(1-z)^{1+{m_{1}+m_{2}\over 2}}&&\quad z\to 1\\ \end{array}\right.$ (51)
@@
$G(r,r^{\prime})={1\over W}\left\{\begin{array}[]{ccc}R_{\rm in}(r^{\prime})R_{\rm up}(r)&{}{}{}{}{}{}{}\hfil&r^{\prime}<r\\ R_{\rm in}(r)R_{\rm up}(r^{\prime})&{}{}{}{}{}{}{}\hfil&r<r^{\prime}\\ \end{array}\right.$ (52)
@@
$R_{\alpha}(y)=e^{-{y\over 2}}\left(1-\frac{x}{y}\right)^{2{-}{x\over 2}}y^{{-}1{-}{x\over 2}}G^{0}_{\alpha}(y)$ (56)
@@
$B^{\rm in,up}_{\alpha^{\prime}}$ $=$ $\sum_{\alpha}c^{\rm in,up}_{\alpha}g_{\alpha}(x)B_{\alpha\alpha^{\prime}}x^{1-\alpha^{\prime}m_{3}}e^{-\frac{1+\alpha^{\prime}}{2}\partial_{m_{3}}{\cal F}_{\rm inst}}$
$D^{\rm in,up}_{\alpha^{\prime}}$ $=$ $\sum_{\alpha}c^{\rm in,up}_{\alpha}F_{\alpha\alpha^{\prime}}h_{\alpha^{\prime}}e^{-{x\over 2}}\,x^{1+m_{3}}$ (58)
@@ 3.2 The incoming solution
$\ {c^{\rm in}_{+}\over c^{\rm in}_{-}}=-{F_{-+}\over F_{++}}=\frac{\Gamma({-}2a)\Gamma\left({\textstyle\frac{1}{2}}{-}m_{1}{+}a\right)\Gamma\left({\textstyle\frac{1}{2}}{-}m_{2}{+}a\right)}{\Gamma(2a)\Gamma\left({\textstyle\frac{1}{2}}{-}m_{1}{-}a\right)\Gamma\left({\textstyle\frac{1}{2}}{-}m_{2}{-}a\right)}$ (63)
@@
$R_{\rm in}(y)\underset{y\to\infty}{\approx}\sum_{\alpha,\alpha^{\prime}}c^{\rm in}_{\alpha}g_{\alpha}B_{\alpha\alpha^{\prime}}e^{\alpha^{\prime}y\over 2}y^{1-\alpha^{\prime}m_{3}}e^{-\frac{1+\alpha^{\prime}}{2}\partial_{m_{3}}{\cal F}_{\rm inst}}$ (64)
@@
$B^{\rm in}_{\alpha^{\prime}}=\sum_{\alpha}c^{\rm in}_{\alpha}g_{\alpha}B_{\alpha\alpha^{\prime}}x^{1-\alpha^{\prime}m_{3}}e^{-\frac{1+\alpha^{\prime}}{2}\partial_{m_{3}}{\cal F}_{\rm inst}}$ (65)
@@
3.3 The upgoing solution
Upgoing boundary conditions correspond to setting $B^{\rm up}_{-}=0$, that according to (58) leads to